OpenStudy (anonymous):

find the limit when x --->0 (cos3x-1)/sin2x

OpenStudy (freckles):

\[\lim_{u \rightarrow 0}\frac{\cos(u)-1}{u}=? \\ \lim_{u \rightarrow 0}\frac{u}{\sin(u)}=?\]

OpenStudy (xapproachesinfinity):

\[\lim_{x \rightarrow 0}\frac{ \cos3x-1 }{ \sin2x }=\lim_{x \rightarrow 0}\frac{ 2x }{ 3x }\frac{ 3x }{2x}\frac{ \cos3x-1}{\sin2x}\]

OpenStudy (xapproachesinfinity):

\[=\lim_{x \rightarrow 0}\frac{ 3 }{ 2}\frac{ \cos3x-1 }{ 3x }\frac{ 2x }{ \sin2x }\] then use the limit of product is the product of limits take the limit of each fraction and you are done!

OpenStudy (xapproachesinfinity):

all the depends on what you know about those two identical limits @freckles has also gave you the hint

OpenStudy (anonymous):

I don't understand the first step

OpenStudy (xapproachesinfinity):

the fist step: i multiplied top and bottom by 2x and also by 3x so i can have something that i know its limit

OpenStudy (xapproachesinfinity):

realize that 6x^2/6x^2 is just 1

OpenStudy (xapproachesinfinity):

if for instance i canceled the 2, 3 and x i would get the initial expression cos3x-1/sin2x, yes?

OpenStudy (xapproachesinfinity):

Do you know the limits @freckles mentioned about or not?

OpenStudy (xapproachesinfinity):

above*